given a 4 percent interest rate, compute the year 6 future value of deposits made in years 1,2,3, and 4 of $1300,$1500,$1500 and $1800 respectively
Answer #1

Year 6 future value of deposits is $ 6,970.61

Working:

Calculation of year 6 value of cash flows:
Year Beginning cash flow Interest Addition in cash flow Ending cash flow
a b=a*4% c d=a+b+c
1 0 0 $ 1,300.00 $ 1,300.00
2 $ 1,300.00 $    52.00 $ 1,500.00 $ 2,852.00
3 $ 2,852.00 $ 114.08 $ 1,500.00 $ 4,466.08
4 $ 4,466.08 $ 178.64 $ 1,800.00 $ 6,444.72
5 $ 6,444.72 $ 257.79 0 $ 6,702.51
6 $ 6,702.51 $ 268.10 0 $ 6,970.61